Operations Report – March 2nd

Crew 223 Operations Report 03/02/2020

SOL: 1

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Sunday generator:
Off at 5:30am
On at 7:00pm SOC was 79%

Monday generator:
Off at 4:30am
On at 9pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149665.40 units

Water (static tank) – 363 gallons

Static to Loft Pump used – yes

Water in GreenHab – 218.3 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: yes
Hours: 149.1
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover: no
Hours: 133.2
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes. Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Run during the engineer EVA for battery
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: The ATV 1’s oil light up.

HabCar used and why, where? Nothing to report.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 4, 7, 8, 9 were used during the EVA. We had troubles with suit number 7 (intermittent fan, same as crew 222 and 221), I tried to fix it but didn’t manage to. I just find an email sent last friday by Scott Davis giving some ideas to fix it, so I will try again tomorrow.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Tried to fix suit 7 without success.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report